Blood is accelerated from rest to 31.0 cm/s in a distance of 1.75 cm by the ventricle of the heart. Determine (a) the acceleration of the blood in cm and (b) the duration of the acceleration
Answer #1

(a) Use the expression -

v² = u² + 2as

Blood is starting from rest. So u = 0

So, v² = 2as

=> a = v²/(2s)

therefore

a = 31²/(2*1.75) = 274.6 cm/s²

(b) Now use the expression -

s = ½(u + v)*t

put u = 0

So, s = ½v*t

=> t = 2s/v

=> t = (2*1.75)/31 = 0.1129 s